40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Oct 4 19:31:17 2011, kiwamu, Update, SUS, PRM and BS oplev laser died oplevs.png
    Reply  Wed Oct 5 11:33:25 2011, steve, Update, SUS, PRM and BS oplev laser replaced 
       Reply  Mon Apr 20 16:42:48 2015, steve, Update, SUS, PRM and BS oplev laser replaced newOplevLaser.png
          Reply  Fri Apr 24 23:40:15 2015, rana, Update, SUS, PRM and BS oplev laser replaced BSOL.pdf
             Reply  Sun Apr 26 15:34:34 2015, Jenne, Update, SUS, PRM oplev centered 
             Reply  Tue Jun 9 10:57:12 2015, Steve, Update, SUS, BS oplev laser tuned 100dBStrend0425.png18dBSdrift.png
Message ID: 5620     Entry time: Wed Oct 5 11:33:25 2011     In reply to: 5618     Reply to this: 11227
Author: steve 
Type: Update 
Category: SUS 
Subject: PRM and BS oplev laser replaced 

 

JDSU 1103P died after 4 years of service. It was replaced with new identical head of 2.9 mW output. The power supply was also changed.

The return spots of 0.04 mW  2.5 mm diameter on qpds are BS  3,700 counts and PRM 4,250 counts.

 

ELOG V3.1.3-